# Exact overlaps for ”all” integrable matrix product states of rational spin chains

### v1

@@ II.1 Definitions for integrable spin chains
In this letter we do not focus on a particular model but we work with a wide class of integrable spin chains. A spin chain is given by a Hamiltonian $H$. The $H$ acts on a Hilbert space $\mathcal{H}$ with a tensor product form $\mathcal{H}=\left(\mathbb{C}^{d}\right)^{J}$ where $d$ is dimension of one-site Hilbert space and $J$ is the length of the spin chain. The model is integrable if there exists a generating function of the commuting charges $\mathcal{T}(u)$ which commute with the Hamiltonian
@@
$\mathcal{T}(u)=\sum_{j=1}^{N}T_{i,i}(u),$ (2)
where $T_{i,j}(u)$ is the entries of the monodromy matrix for $i,j=1,\dots,N$. We also introduce the matrix notation in the auxiliary space $T(u)=\sum_{i,j}e_{i,j}\otimes T_{i,j}(u)\in\mathrm{End}(\mathbb{C}^{N}\otimes\mathcal{H})$, where $e_{i,j}$ are the unit matrices of $\mathbb{C}^{N}$ which have the components $\left(e_{i,j}\right)_{a,b}=\delta_{i,a}\delta_{j,b}$. The monodromy matrix satisfies the famous $RTT$-relation
@@ II.2 Integrable matrix product states
where $\psi_{i,j}^{(k)}\in\mathrm{End}(\mathbb{C}^{d_{B}})$ for $i,j=1,\dots,d$, $k=1,\dots,J/2$ and $d_{B}$ is the boundary (or bond) dimension. The $\psi_{i,j}^{(k)}$ matrices can be different (i.e. $\psi_{i,j}^{(k)}\neq\psi_{i,j}^{(l)}$) for inhomogeneous chain but they are given by the same functions as $\psi_{i,j}^{(k)}\equiv\psi_{i,j}(\theta_{k})$, where $\theta_{k}$-s are the inhomogeneities. For homogenous chains it can happen that $\psi_{i,j}^{(k)}=\omega_{j}\omega_{i}$ for which the MPS can be also defined for even or odd length:
@@
where the coefficients $K_{i,j}^{\alpha,\beta}(u)\in\mathbb{C}$ define the K-matrix $\mathbf{K}_{i,j}(u)=\sum_{\alpha,\beta}K_{i,j}^{\alpha,\beta}(u)e_{\alpha,\beta}^{B}\in\mathrm{End}(\mathbb{C}^{d_{B}})$, $\mathbf{K}(u)=\sum_{i,j}e_{i,j}\otimes\mathbf{K}_{i,j}(u)\in\mathrm{End}(\mathbb{C}^{N}\otimes\mathbb{C}^{d_{B}})$. We used the unit matrices $(e_{\alpha,\beta}^{B})_{a,b}=\delta_{\alpha,a}\delta_{\beta,b}$ of the boundary space $\mathbb{C}^{d_{B}}$. We also define the notation $\langle\Psi|\in\mathcal{H}^{*}\otimes\mathrm{End}(\mathbb{C}^{d_{B}})$ which is a covector in the Hilbert space and a matrix in the boundary space: $\langle\Psi|=\sum_{\alpha,\beta}\langle\Psi_{\alpha,\beta}|\otimes e_{\alpha,\beta}^{B}$. Using these matrix notations in the boundary space the $KT$-relation has the following compact form
@@ III.1 Pair structures
Now we continue with the overlaps between the integrable MPS-s and the on-shell Bethe states. The integrability condition (17) is equivalent to non-trivial selection rules for the on-shell overlaps, namely, the on-shell overlap is non-vanishing only when the Bethe roots have pair structures, i.e. $\bar{u}^{j}=-\bar{u}^{\rho(j)}$ for $j=1,\dots,n$. The $\rho$ is an involution of the Dynkin-diagram. If $\rho$ is trivial, i.e., $\rho(j)=j$ for every $j$ then we have chiral pair structure. The pair structure is achiral for non-trivial $\rho$. We have two possibilities for non-trivial $\rho$: for $\mathfrak{gl}_{N}$ $\rho(j)=N-j$; and for $\mathfrak{o}_{2n}$ $\rho(n-1)=n$, $\rho(n)=n-1$ and $\rho(j)=j$ for $j<n-1$. For $\mathfrak{o}_{2n+1}$ and $\mathfrak{so}_{2n}$ there is no non-trivial involution therefore only the chiral pair structure is possible. Every twisted Yangian have a specific pair structure which is summarized in table 1. Integrable MPS with a given chirality can only exist for spin chains where the following conditions are satisfied $\alpha_{j}(u)\alpha_{\rho(j)}(-u)=1$ for $j=1,\dots,n$.
@@ III.2 Overlap formula
The prefactor (blue) is the MPS dependent part. We introduced shorthand notation ~$\tilde{\mathcal{F}}_{k}^{(j)}(\bar{v})=\prod_{v_{j}\in\bar{v}}\tilde{\mathcal{F}}_{k}^{(j)}(v)$ where ~$\tilde{\mathcal{F}}_{k}^{(j)}(u)$ are single-variable functions for $j=1,\dots,n_{+}$ and $k=1,\dots,d_{B}$. These functions ~$\tilde{\mathcal{F}}_{k}^{(j)}(u)$ depend only on the $K$-matrix, but not on the concrete representation of the Hilbert space. For every twisted Yangian $Y(\mathfrak{g},\mathfrak{h})$ there is a commuting subalgebra which are generated by $\mathbf{F}^{(j)}(u)$ for $j=1,\dots,n_{+}$. The $F$-operators $\mathbf{F}^{(j)}(u)$ can be expressed by the components of the $K$-matrix. Since these are commuting operators they can be diagonalized simultaneously and the eigenvalues give ~$\tilde{\mathcal{F}}_{k}^{(j)}(u)$.
@@ III.3 Example $Y(\mathfrak{gl}_{N},\mathfrak{o}_{N})$
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\mathcal{F}_{j}^{(k)}(-iu-k/2)\sqrt{\frac{u^{2}}{u^{2}+1/4}}.$ (26)
@@ 3. Identify the reflection algebra
Now we have to find the twisted Yangian $Y(\mathfrak{g},\mathfrak{h})$ algebra corresponding to our K-matrix. For this we have to identify the Lie-subalgebra $\mathfrak{h}$ from the series expansion of the $K$-matrix (20). The knowledge of twisted Yangian $Y(\mathfrak{g},\mathfrak{h})$ already selects the pair of structures which defines the Gaudin determinants (see Supplemental Materials).
@@
where $\hat{g}=g$ except for the crossed $\mathfrak{gl}_{N}$ when $\hat{g}=\left(g^{t}\right)^{-1}$. The $\Delta(g)$ is usual co-product corresponding to the Hilbert space. The $G$-action preserves the $KT$-relation. Our task is to find a group element $g(\varphi)=\exp(\varphi x)$ ($x\in\mathfrak{g}$,$\varphi\in\mathbb{C}$) for which the definition of the $F$-operators are well defined. The further steps are performed on the ”rotated” K-matrix $\mathbf{K}^{g(\varphi)}$ (if the rotation is necessary), and we can take the $\varphi\to 0$ limit in the final formula.
@@ 5. Calculation of the F-operators
Since we know the $K$-matrix and the associated reflection algebra $Y(\mathfrak{g},\mathfrak{h})$, all we have to do is to calculate the $F$-operators based on section of $Y(\mathfrak{g},\mathfrak{h})$ in the Supplemental Materials. Their eigenvalues define the functions ~$\tilde{\mathcal{F}}_{k}^{(j)}(u)$.
@@ ACKNOWLEDGMENTS
Supplemental Materials: Wrapping corrections for long range spin chains
@@
    $\begin{split}e^{i\phi_{k}^{a}}&:=\alpha_{a}(u_{k}^{a})\prod_{l=1}^{r_{a}}\frac{u_{k}^{a}-u_{l}^{a}-i}{u_{k}^{a}-u_{l}^{a}+i}\prod_{l=1}^{r_{a-1}}\frac{u_{k}^{a}-u_{l}^{a-1}+\frac{i}{2}}{u_{k}^{a}-u_{l}^{a-1}-\frac{i}{2}}\prod_{l=1}^{r_{a+1}}\frac{u_{k}^{a}-u_{l}^{a+1}+\frac{i}{2}}{u_{k}^{a}-u_{l}^{a+1}-\frac{i}{2}}=-1,\quad a=1,\dots,n-3,\\ e^{i\phi_{k}^{n-2}}&:=\alpha_{n-2}(u_{k}^{n-2})\prod_{l=1}^{r_{n-2}}\frac{u_{k}^{n-2}-u_{l}^{n-2}-i}{u_{k}^{n-2}-u_{l}^{n-2}+i}\prod_{l=1}^{r_{n-3}}\frac{u_{k}^{n-2}-u_{l}^{n-3}+\frac{i}{2}}{u_{k}^{n-2}-u_{l}^{n-3}-\frac{i}{2}}\times\\ &\times\prod_{l=1}^{r_{n-1}}\frac{u_{k}^{n-2}-u_{l}^{n-1}+i/2}{u_{k}^{n-2}-u_{l}^{n-1}-i/2}\prod_{l=1}^{r_{n}}\frac{u_{k}^{n-2}-u_{l}^{n}+i/2}{u_{k}^{n-2}-u_{l}^{n}-i/2}=-1,\\ e^{i\phi_{k}^{n-1}}&:=\alpha_{n-1}(u_{k}^{n-1})\prod_{l=1}^{r_{n-1}}\frac{u_{k}^{n-1}-u_{l}^{n-1}-i}{u_{k}^{n-1}-u_{l}^{n-1}+i}\prod_{l=1}^{r_{n-2}}\frac{u_{k}^{n-1}-u_{l}^{n-2}+\frac{i}{2}}{u_{k}^{n-1}-u_{l}^{n-2}-\frac{i}{2}}=-1,\\ e^{i\phi_{k}^{n}}&:=\alpha_{n}(u_{k}^{n})\prod_{l=1}^{r_{n}}\frac{u_{k}^{n}-u_{l}^{n}-i}{u_{k}^{n}-u_{l}^{n}+i}\prod_{l=1}^{r_{n-1}}\frac{u_{k}^{n}-u_{l}^{n-2}+i/2}{u_{k}^{n}-u_{l}^{n-2}-i/2}=-1.\end{split}$ (S-5)
@@
For chiral pair structures $\bar{u}^{s}=-\bar{u}^{s}$ for $s=1,\dots,n$. We define the quantum numbers $n_{+}=n$ and $r_{s}^{+}=r_{s}/2$ for $s=1,\dots,n$. We introduce a new notations for the Bethe roots $u_{k}^{+,s}\equiv u_{k}^{s}$, $u_{k}^{-,s}\equiv u_{r_{s}+1-k}^{s}$ where $s=1,\dots,n_{+}$ and $k=1,\dots,r_{s}^{+}$. We also define the following sets $\bar{u}^{\pm,s}=\left\{u_{k}^{\pm,s}\right\}_{k=1}^{r_{s}^{+}}$ for $s=1,\dots,n_{+}$ and $\bar{u}^{\pm}=\left\{\bar{u}^{\pm,1},\bar{u}^{\pm,2},\dots,\bar{u}^{\pm,n_{+}}\right\}$. In the pair structure limit $u_{k}^{-,s}=-u_{k}^{+,s}$ for $s=1,\dots,n_{+}$ and $k=1,\dots,r_{s}^{+}$. Chiral integrable MPS exists only for the spin chains where the following conditions are satisfied: $\alpha_{s}(u)\alpha_{s}(-u)=1$ for $s=1,\dots,n$.
@@
  • For $\mathfrak{o}_{2n}$, the achiral pair structure is $\bar{u}^{1}=-\bar{u}^{1}$, $\bar{u}^{2}=-\bar{u}^{2}$, $\dots$,$\bar{u}^{n-2}=-\bar{u}^{n-2}$ and $\bar{u}^{n}=-\bar{u}^{n-1}$. We can see that the nodes $\bar{u}^{s}$ for $s=1,\dots,n-2$ are chiral. We define the quantum numbers $n_{+}=n-1$, $r_{s}^{+}=r_{s}/2$ for $s=1,\dots,n-2$ and $r_{n_{+}}^{+}=r_{n-1}=r_{n}$. We introduce a new notations for Bethe roots $u_{k}^{+,s}\equiv u_{k}^{s}$, $u_{k}^{-,s}\equiv u_{r_{s}+1-k}^{s}$ where $s=1,\dots,n_{+}-1$, $k=1,\dots,r_{s}^{+}$ and $u_{k}^{+,n_{+}}\equiv u_{k}^{n-1}$, $u_{k}^{-,n_{+}}\equiv u_{k}^{n}$ where $k=1,\dots,r_{n-1}^{+}$. Achiral integrable MPS-s of $\mathfrak{o}_{2n}$ symmetric spin chains exist only when the following conditions are satisfied: $\alpha_{s}(u)\alpha_{s}(-u)=1$ for $s=1,\dots,n-2$ and $\alpha_{n-1}(u)\alpha_{n}(-u)=1$.
@@
The shorthand notation for the Bethe Ansatz equations is
$\phi_{k}^{s}=0,$ (S-6)
for $s=1,\dots,n$ and $k=1,\dots,r_{s}$. The expressions $\phi_{k}^{s}$ are depend on the Bethe roots and they are defined above. These can be grouped according to the pair structure in the same way as Bethe roots: $\bar{\phi}^{\pm,s}=\left\{\phi_{k}^{\pm,s}\right\}_{k=1}^{r_{s}^{+}}$ where $s=1,\dots,n_{+}$ and $\bar{\phi}^{\pm}=\left\{\bar{\phi}^{\pm,1},\bar{\phi}^{\pm,2},\dots,\bar{\phi}^{\pm,n_{+}}\right\}$. We can define the components of the Gaudin matrix
$G_{k,l}^{\mu,\nu}=\frac{\partial\phi_{k}^{\mu}}{\partial u_{l}^{\nu}}$ (S-7)
The Gaudin matrices are have $\sum_{s=1}^{n}r_{s}$ rows and columns and the rows are labeled by $\nu,l$ and the columns are labeled by $\mu,k$. According to the decompositions $\bar{u}=\bar{u}^{+}\cup\bar{u}^{-}$ and $\bar{\phi}=\bar{\phi}^{+}\cup\bar{\phi}^{-}$ the Gaudin matrix have the following block form
$G=\left(\begin{array}[]{cc}A^{++}&A^{+-}\\ A^{-+}&A^{--}\end{array}\right),$ (S-8)
@@
$\left(A^{\sigma_{1}\sigma_{2}}\right)_{k,l}^{\mu,\nu}=\frac{\partial\phi_{k}^{\sigma_{1},\mu}}{\partial u_{l}^{\sigma_{2},\nu}},$ (S-9)
@@
$\alpha_{s}(u)=1+\frac{i}{u}d_{s}+\mathcal{O}(u^{-2}),$ (S-11)
@@
$Y(\mathfrak{gl}_{2n},\mathfrak{sp}_{2n})$ (chiral)
For the $Y(\mathfrak{gl}_{2n},\mathfrak{sp}_{2n})$ the overlap is non-vanishing only when
$r_{2a-1}=\frac{d_{2a-1}+r_{2a-2}+r_{2a}}{2},$ (S-12)
for $a=1,\dots,n$. We can see that the pseudo-vacuum overlap $\mathbf{B}=\langle\Psi|0\rangle$ is non-vanishing only when $d_{2a-1}=0$ for $a=1,\dots,n$.
@@
$\Lambda_{j}=\begin{cases}\sum_{k=j}^{n-1}d_{k}+\frac{1}{2}d_{n},&N=2n\\ \sum_{k=j}^{n}d_{k},&N=2n+1\end{cases}$ (S-14)
@@
$\begin{split}r_{2a-1}&=\frac{r_{2a-2}+r_{2a}+d_{2a-1}}{2},\quad\text{for }a=1,\dots,m,\\ r_{j}&=r_{2m}+\sum_{k=2m+1}^{j}\Lambda_{k},\quad j=2m+1,\dots,n-1,\\ r_{n}&=\frac{1}{2}r_{2m}+\frac{1}{2}\sum_{k=2m+1}^{n}\Lambda_{n},\end{split}$ (S-17)
@@
$\begin{split}r_{j}&=r_{M}+\sum_{k=M+1}^{j}\Lambda_{k},\quad j=M+1,\dots,n-2,\\ r_{n-1}&=\frac{1}{2}r_{M}+\frac{1}{2}\sum_{k=M+1}^{n-1}\Lambda_{k}-\frac{1}{2}\Lambda_{n},\\ r_{n}&=\frac{1}{2}r_{M}+\frac{1}{2}\sum_{k=M+1}^{n-1}\Lambda_{k}+\frac{1}{2}\Lambda_{n}.\end{split}$ (S-19)
@@
$\begin{split}\Lambda_{j}&=\sum_{k=j}^{n-2}d_{k}+\frac{1}{2}d_{n-1}+\frac{1}{2}d_{n},\quad j=M+1,\dots,n-1\\ \Lambda_{n}&=\frac{1}{2}d_{n}-\frac{1}{2}d_{n-1}\end{split}$ (S-20)
@@
$\begin{split}r_{2a-1}&=\frac{r_{2a-2}+r_{2a}+d_{2a-1}}{2},\quad\text{for }a=1,\dots,\frac{n-3}{2},\\ r_{n-2}&=\frac{r_{n-3}+2r_{n-1}+d_{n-2}}{2}.\end{split}$ (S-22)

### v2

@@ II.1 Definitions for integrable spin chains
In this letter we do not focus on a particular model but we work with a wide class of integrable spin chains. A spin chain is given by a Hamiltonian $H$. The $H$ acts on a Hilbert space $\mathcal{H}$ with a tensor product form $\mathcal{H}=\left(\mathbb{C}^{d}\right)^{J}$ where $d$ is the dimension of the one-site Hilbert space and $J$ is the length of the spin chain. The model is integrable if there exists a generating function $\mathcal{T}(u)$ of the commuting charges which commute with the Hamiltonian
@@
$\mathcal{T}(u)=\sum_{j=1}^{N}T_{j,j}(u),$ (2)
where $T_{i,j}(u)$ are the entries of the monodromy matrix for $i,j=1,\dots,N$. We also introduce the matrix notation in the auxiliary space $T(u)=\sum_{i,j}e_{i,j}\otimes T_{i,j}(u)\in\mathrm{End}(\mathbb{C}^{N}\otimes\mathcal{H})$, where $e_{i,j}$ are the unit matrices of $\mathbb{C}^{N}$ which have the components $\left(e_{i,j}\right)_{a,b}=\delta_{i,a}\delta_{j,b}$. The monodromy matrix satisfies the famous $RTT$-relation
@@ II.2 Integrable matrix product states
where $\psi_{i,j}^{(k)}\in\mathrm{End}(\mathbb{C}^{d_{B}})$ for $i,j=1,\dots,d$, $k=1,\dots,J/2$ and $d_{B}$ is the boundary (or bond) dimension. The $\psi_{i,j}^{(k)}$ matrices can be different (i.e. $\psi_{i,j}^{(k)}\neq\psi_{i,j}^{(l)}$) for inhomogeneous chains but they are given by the same functions as $\psi_{i,j}^{(k)}\equiv\psi_{i,j}(\theta_{k})$, where $\theta_{k}$-s are the inhomogeneities. For homogenous chains it can happen that $\psi_{i,j}^{(k)}=\omega_{j}\omega_{i}$ for which the MPS can be also defined for even or odd length:
@@
where the coefficients $K_{i,j}^{\alpha,\beta}(u)\in\mathbb{C}$ define the K-matrix $\mathbf{K}_{i,j}(u)=\sum_{\alpha,\beta}K_{i,j}^{\alpha,\beta}(u)e_{\alpha,\beta}^{B}\in\mathrm{End}(\mathbb{C}^{d_{B}})$, $\mathbf{K}(u)=\sum_{i,j}e_{i,j}\otimes\mathbf{K}_{i,j}(u)\in\mathrm{End}(\mathbb{C}^{N}\otimes\mathbb{C}^{d_{B}})$. We used the elementary matrices $(e_{\alpha,\beta}^{B})_{a,b}=\delta_{\alpha,a}\delta_{\beta,b}$ of the boundary space $\mathbb{C}^{d_{B}}$. We also define the notation $\langle\Psi|\in\mathcal{H}^{*}\otimes\mathrm{End}(\mathbb{C}^{d_{B}})$ which is a covector in the Hilbert space and a matrix in the boundary space: $\langle\Psi|=\sum_{\alpha,\beta}\langle\Psi_{\alpha,\beta}|\otimes e_{\alpha,\beta}^{B}$. Using these matrix notations in the boundary space the $KT$-relation has the following compact form
@@ III.1 Pair structures
Now we continue with the overlaps between the integrable MPS-s and the on-shell Bethe states. The integrability condition (17) is equivalent to non-trivial selection rules for the on-shell overlaps, namely, the on-shell overlap is non-vanishing only when the Bethe roots have pair structures, i.e. $\bar{u}^{j}=-\bar{u}^{\rho(j)}$ for $j=1,\dots,n$. The $\rho$ is an involution of the Dynkin-diagram. If $\rho$ is trivial, i.e., $\rho(j)=j$ for every $j$ then we have chiral pair structure. The pair structure is achiral for non-trivial $\rho$. We have two possibilities for non-trivial $\rho$: for $\mathfrak{gl}_{N}$ $\rho(j)=N-j$; and for $\mathfrak{o}_{2n}$ $\rho(n-1)=n$, $\rho(n)=n-1$ and $\rho(j)=j$ for $j<n-1$. For $\mathfrak{o}_{2n+1}$ and $\mathfrak{so}_{2n}$ there is no non-trivial involution therefore only the chiral pair structure is possible. Every twisted Yangian has a specific pair structure which is summarized in table 1. Integrable MPS with a given chirality can only exist for spin chains where the following conditions are satisfied $\alpha_{j}(u)\alpha_{\rho(j)}(-u)=1$ for $j=1,\dots,n$.
@@ III.2 Overlap formula
The prefactor (blue) is the MPS dependent part. We introduced shorthand notation ~$\tilde{\mathcal{F}}_{k}^{(j)}(\bar{v})=\prod_{v_{l}\in\bar{v}}\tilde{\mathcal{F}}_{k}^{(j)}(v_{l})$ where ~$\tilde{\mathcal{F}}_{k}^{(j)}(u)$-s are single-variable functions for $j=1,\dots,n_{+}$ and $k=1,\dots,d_{B}$. These functions ~$\tilde{\mathcal{F}}_{k}^{(j)}(u)$ depend only on the $K$-matrix, but not on the concrete representation of the Hilbert space. For every twisted Yangian $Y(\mathfrak{g},\mathfrak{h})$ there is a commuting subalgebra which are generated by $\mathbf{F}^{(j)}(u)$ for $j=1,\dots,n_{+}$. The $F$-operators $\mathbf{F}^{(j)}(u)$ can be expressed by the components of the $K$-matrix. Since these are commuting operators they can be diagonalized simultaneously and the eigenvalues give ~$\tilde{\mathcal{F}}_{k}^{(j)}(u)$.
@@ III.3 Example $Y(\mathfrak{gl}_{N},\mathfrak{o}_{N})$
~$\tilde{\mathcal{F}}_{j}^{(k)}(u)=\mathcal{F}_{j}^{(k)}(iu-k/2)\sqrt{\frac{u^{2}}{u^{2}+1/4}}.$ (26)
@@ 3. Identify the reflection algebra
Now we have to find the twisted Yangian $Y(\mathfrak{g},\mathfrak{h})$ algebra corresponding to our K-matrix. For this we have to identify the Lie-subalgebra $\mathfrak{h}$ from the series expansion of the $K$-matrix (20). The knowledge of twisted Yangian $Y(\mathfrak{g},\mathfrak{h})$ already selects the pair structure which defines the Gaudin determinants (see Supplemental Materials).
@@
where $\hat{g}=g$ except for the crossed $\mathfrak{gl}_{N}$ when $\hat{g}=\left(g^{t}\right)^{-1}$. The $\Delta(g)$ is the usual co-product corresponding to the Hilbert space. The $G$-action preserves the $KT$-relation. Our task is to find a group element $g(\varphi)=\exp(\varphi x)$ ($x\in\mathfrak{g}$,$\varphi\in\mathbb{C}$) for which the definition of the $F$-operators are well defined. The further steps are performed on the ”rotated” K-matrix $\mathbf{K}^{g(\varphi)}$ (if the rotation is necessary), and we can take the $\varphi\to 0$ limit in the final formula.
@@ 5. Calculation of the F-operators
Since we know the $K$-matrix and the associated reflection algebra $Y(\mathfrak{g},\mathfrak{h})$, all we have to do is to calculate the $F$-operators based on the section of $Y(\mathfrak{g},\mathfrak{h})$ in the Supplemental Materials. Their eigenvalues define the functions ~$\tilde{\mathcal{F}}_{k}^{(j)}(u)$.
@@ Acknowledgments
Supplemental Materials: Exact overlaps for ”all” integrable matrix product states of rational spin chains
@@
    $\begin{split}e^{i\phi_{k}^{a}}&:=\alpha_{a}(u_{k}^{a})\prod_{l=1}^{r_{a}}\frac{u_{k}^{a}-u_{l}^{a}-i}{u_{k}^{a}-u_{l}^{a}+i}\prod_{l=1}^{r_{a-1}}\frac{u_{k}^{a}-u_{l}^{a-1}+\frac{i}{2}}{u_{k}^{a}-u_{l}^{a-1}-\frac{i}{2}}\prod_{l=1}^{r_{a+1}}\frac{u_{k}^{a}-u_{l}^{a+1}+\frac{i}{2}}{u_{k}^{a}-u_{l}^{a+1}-\frac{i}{2}}=-1,\quad a=1,\dots,n-3,\\ e^{i\phi_{k}^{n-2}}&:=\alpha_{n-2}(u_{k}^{n-2})\prod_{l=1}^{r_{n-2}}\frac{u_{k}^{n-2}-u_{l}^{n-2}-i}{u_{k}^{n-2}-u_{l}^{n-2}+i}\prod_{l=1}^{r_{n-3}}\frac{u_{k}^{n-2}-u_{l}^{n-3}+\frac{i}{2}}{u_{k}^{n-2}-u_{l}^{n-3}-\frac{i}{2}}\times\\ &\times\prod_{l=1}^{r_{n-1}}\frac{u_{k}^{n-2}-u_{l}^{n-1}+\frac{i}{2}}{u_{k}^{n-2}-u_{l}^{n-1}-\frac{i}{2}}\prod_{l=1}^{r_{n}}\frac{u_{k}^{n-2}-u_{l}^{n}+\frac{i}{2}}{u_{k}^{n-2}-u_{l}^{n}-\frac{i}{2}}=-1,\\ e^{i\phi_{k}^{n-1}}&:=\alpha_{n-1}(u_{k}^{n-1})\prod_{l=1}^{r_{n-1}}\frac{u_{k}^{n-1}-u_{l}^{n-1}-i}{u_{k}^{n-1}-u_{l}^{n-1}+i}\prod_{l=1}^{r_{n-2}}\frac{u_{k}^{n-1}-u_{l}^{n-2}+\frac{i}{2}}{u_{k}^{n-1}-u_{l}^{n-2}-\frac{i}{2}}=-1,\\ e^{i\phi_{k}^{n}}&:=\alpha_{n}(u_{k}^{n})\prod_{l=1}^{r_{n}}\frac{u_{k}^{n}-u_{l}^{n}-i}{u_{k}^{n}-u_{l}^{n}+i}\prod_{l=1}^{r_{n-2}}\frac{u_{k}^{n}-u_{l}^{n-2}+\frac{i}{2}}{u_{k}^{n}-u_{l}^{n-2}-\frac{i}{2}}=-1.\end{split}$ (S-5)
@@
For chiral pair structures $\bar{u}^{s}=-\bar{u}^{s}$ for $s=1,\dots,n$. We define the quantum numbers $n_{+}=n$ and $r_{s}^{+}=r_{s}/2$ for $s=1,\dots,n$. We introduce new notations for the Bethe roots $u_{k}^{+,s}\equiv u_{k}^{s}$, $u_{k}^{-,s}\equiv u_{r_{s}+1-k}^{s}$ where $s=1,\dots,n_{+}$ and $k=1,\dots,r_{s}^{+}$. We also define the following sets $\bar{u}^{\pm,s}=\left\{u_{k}^{\pm,s}\right\}_{k=1}^{r_{s}^{+}}$ for $s=1,\dots,n_{+}$ and $\bar{u}^{\pm}=\left\{\bar{u}^{\pm,1},\bar{u}^{\pm,2},\dots,\bar{u}^{\pm,n_{+}}\right\}$. In the pair structure limit $u_{k}^{-,s}=-u_{k}^{+,s}$ for $s=1,\dots,n_{+}$ and $k=1,\dots,r_{s}^{+}$. Chiral integrable MPS exists only for the spin chains where the following conditions are satisfied: $\alpha_{s}(u)\alpha_{s}(-u)=1$ for $s=1,\dots,n$.
@@
  • For $\mathfrak{o}_{2n}$, the achiral pair structure is $\bar{u}^{1}=-\bar{u}^{1}$, $\bar{u}^{2}=-\bar{u}^{2}$, $\dots$,$\bar{u}^{n-2}=-\bar{u}^{n-2}$ and $\bar{u}^{n}=-\bar{u}^{n-1}$. We can see that the nodes $\bar{u}^{s}$ for $s=1,\dots,n-2$ are chiral. We define the quantum numbers $n_{+}=n-1$, $r_{s}^{+}=r_{s}/2$ for $s=1,\dots,n-2$ and $r_{n_{+}}^{+}=r_{n-1}=r_{n}$. We introduce new notations for Bethe roots $u_{k}^{+,s}\equiv u_{k}^{s}$, $u_{k}^{-,s}\equiv u_{r_{s}+1-k}^{s}$ where $s=1,\dots,n_{+}-1$, $k=1,\dots,r_{s}^{+}$ and $u_{k}^{+,n_{+}}\equiv u_{k}^{n-1}$, $u_{k}^{-,n_{+}}\equiv u_{k}^{n}$ where $k=1,\dots,r_{n-1}^{+}$. Achiral integrable MPS-s of $\mathfrak{o}_{2n}$ symmetric spin chains exist only when the following conditions are satisfied: $\alpha_{s}(u)\alpha_{s}(-u)=1$ for $s=1,\dots,n-2$ and $\alpha_{n-1}(u)\alpha_{n}(-u)=1$.
@@
The expressions $\phi_{k}^{s}$ are depend on the Bethe roots and they are defined above. These can be grouped according to the pair structure in the same way as the Bethe roots: $\bar{\phi}^{\pm,s}=\left\{\phi_{k}^{\pm,s}\right\}_{k=1}^{r_{s}^{+}}$ where $s=1,\dots,n_{+}$ and $\bar{\phi}^{\pm}=\left\{\bar{\phi}^{\pm,1},\bar{\phi}^{\pm,2},\dots,\bar{\phi}^{\pm,n_{+}}\right\}$. We can define the components of the Gaudin matrix
$G_{k,l}^{\mu,\nu}=\frac{\partial\phi_{k}^{\mu}}{\partial u_{l}^{\nu}}.$ (S-6)
The Gaudin matrices have $\sum_{s=1}^{n}r_{s}$ rows and columns and the rows are labeled by $\nu,l$ and the columns are labeled by $\mu,k$. According to the decompositions $\bar{u}=\bar{u}^{+}\cup\bar{u}^{-}$ and $\bar{\phi}=\bar{\phi}^{+}\cup\bar{\phi}^{-}$ the Gaudin matrix has the following block form
$G=\left(\begin{array}[]{cc}A^{++}&A^{+-}\\ A^{-+}&A^{--}\end{array}\right),$ (S-7)
@@
$\left(A^{\sigma_{1}\sigma_{2}}\right)_{k,l}^{\mu,\nu}=\frac{\partial\phi_{k}^{\sigma_{1},\mu}}{\partial u_{l}^{\sigma_{2},\nu}},$ (S-8)
@@
$\alpha_{s}(u)=1+\frac{i}{u}d_{s}+\mathcal{O}(u^{-2}),$ (S-10)
@@
$Y(\mathfrak{gl}_{2k},\mathfrak{sp}_{2k})$ (chiral)
For the $Y(\mathfrak{gl}_{2k},\mathfrak{sp}_{2k})$ the overlap is non-vanishing only when the numbers $r_{2a}$ are even for $a=1,\dots,k-1$, and
$r_{2a-1}=\frac{d_{2a-1}+r_{2a-2}+r_{2a}}{2},$ (S-11)
for $a=1,\dots,k$. We can see that the pseudo-vacuum overlap $\mathbf{B}=\langle\Psi|0\rangle$ is non-vanishing only when $d_{2a-1}=0$ for $a=1,\dots,k$.
@@
$\Lambda_{j}=\begin{cases}\sum_{k=j}^{n-1}d_{k}+\frac{1}{2}d_{n},&N=2n,\\ \sum_{k=j}^{n}d_{k},&N=2n+1,\end{cases}$ (S-13)
@@
$\begin{split}r_{2a-1}&=\frac{r_{2a-2}+r_{2a}+d_{2a-1}}{2},\quad\text{for }a=1,\dots,m,\\ r_{j}&=r_{2m}+\sum_{k=2m+1}^{j}\Lambda_{k},\quad j=2m+1,\dots,n-1,\\ r_{n}&=\frac{1}{2}r_{2m}+\frac{1}{2}\sum_{k=2m+1}^{n}\Lambda_{k},\end{split}$ (S-16)
@@
$\begin{split}r_{j}&=r_{M}+\sum_{k=M+1}^{j}\Lambda_{k},\quad j=M+1,\dots,n-2,\\ r_{n-1}&=\frac{1}{2}r_{M}+\frac{1}{2}\sum_{k=M+1}^{n-1}\Lambda_{k}-\frac{1}{2}\Lambda_{n},\\ r_{n}&=\frac{1}{2}r_{M}+\frac{1}{2}\sum_{k=M+1}^{n-1}\Lambda_{k}+\frac{1}{2}\Lambda_{n},\end{split}$ (S-18)
@@
$\begin{split}\Lambda_{j}&=\sum_{k=j}^{n-2}d_{k}+\frac{1}{2}d_{n-1}+\frac{1}{2}d_{n},\quad j=M+1,\dots,n-1,\\ \Lambda_{n}&=\frac{1}{2}d_{n}-\frac{1}{2}d_{n-1}.\end{split}$ (S-19)
@@
$\begin{split}r_{2a-1}&=\frac{r_{2a-2}+r_{2a}+d_{2a-1}}{2},\quad\text{for }a=1,\dots,\frac{n-3}{2},\\ r_{n-2}&=\frac{r_{n-3}+2r_{n-1}+d_{n-2}}{2}.\end{split}$ (S-21)